Part A Calculate the grams of solute in 25.0 mL of a 2.50%m/v solution of fructose. IVO AND A 0 0 ? Submit Request Answer
Answer #1

Solution:-

​​​​​​% m/v = [mass of solute (in grams) / volume of solution (in mL)] × 100%

2.5 = (mass of solute / 25) × 100

(2.5/100) = mass of solute / 25

or, mass of solute = (2.5 / 100) × 25 = 0.625 g

Hence, mass of 25.0 mL of 2.50% m/v solution of fructose is = 0.625 g

---------------------------------------------------------------------------